OpenStudy (anonymous):

what is the vertex form of the equation? y=x^2 +4x-3

OpenStudy (zale101):

y = a(x – h)^2 + k, where (h, k) is the vertex

OpenStudy (anonymous):

so 4,3 would be the vertex?

OpenStudy (anonymous):

y=(x-2)^2-7 y=(x+2)^2-7 y=(x-2)^2+7 y=(x+2)^2+7 are the possible answers

OpenStudy (zale101):

a = 1, b = 4, and c = -3 plug n the quadratic formula